OpenStudy (anonymous):

Given the equation )x+3)^2+y^2=16 , (a) find the center and radius; (b) graph; (c) find the intercepts, if any. Please show all of your work

OpenStudy (anonymous):

 At any point \[r^2=x^2+y^2\] Because of Pythagoras. The only thing that would satisfy this equation is a circle. So, \[r^2=16 \therefore r=4\] Replacing x with (x+3) just complicates it a little. Just move the centre 3 to the left, negative end. So the intercepts are (on the x-axis) -3 (the centre)+4 and -3-4 Which are -7 and 1, which is you insert into the first equation gives y=0